# Thucy: An LLM-based Multi-Agent System for Claim Verification across Relational Databases

### v1

@@ 1 Introduction
In the Annual Report released last year by the Seattle City Attorney’s Office (2024), we read the following:
@@
It turns out, the City of Seattle publicly provides an official crime dataset (Seattle Police Department 2025b)—with all crimes from 2008 until now—that is structured, detailed, and updated almost daily. In principle, that is all you would need to verify such a claim. In practice, though, very few people ever try. Most will simply take the statement at face value and move on, keeping the comforting thought that “Seattle is safer now” somewhere in the back of their mind.
@@
We draw inspiration from the work of Thucy(dides), the Athenian historian (460–400 BC) who wrote the History of the Peloponnesian War between Sparta and Athens. “Thucydides has been dubbed the father of scientific history by those who accept his claims to have applied strict standards of impartiality and evidence-gathering” (Wikipedia 2025).
@@ 2 Architecture
BINDER (Cheng et al. 2023)
DATER (Ye et al. 2023)
CoTable (Wang et al. 2024)
ReAcTable (Zhang et al. 2024)
AutoTQA (Zhu et al. 2024)
POS (Nguyen et al. 2025)
@@
To enable this flexibility, we must address a fundamental challenge: LLMs, no matter how capable, are inherently disconnected from external data sources—they can only operate in isolation, with no way to interact with databases. Agents bridge this gap by using tools. A tool acts as an interface to external capabilities, allowing agents to interact with, perceive, and affect their environment. In general, tools can include capabilities that perform mathematical calculations, or read files from disk, or query a database. Each agent has a fixed collection of such tools. At runtime, the agent [^1] More precisely, it is the LLM that makes this decision, though we often use “agent” and “LLM” interchangeably in such contexts autonomously decides which tool to invoke, how to call it, and when to use it; the tool’s output is then fed back into the reasoning loop (Yao et al. 2023). This interactive feedback cycle between reasoning, action, and observations forms the backbone of modern agentic AI.
@@
The Model Context Protocol (MCP), introduced by Anthropic (2024), standardizes how AI applications connect to different data sources, effectively eliminating the need for custom connections for each new AI model and external system allowing us to direct our energy elsewhere—away from repetitive boilerplate code. MCP simplifies and streamlines the tool building process, and has already been adopted by industry (Mehrotra 2025; Gonzales and Murching 2025; Ganguly 2025; Agarwal et al. 2025).
@@
We use Google’s MCP Toolbox for Databases (Buvaraghan and Egan 2025), a framework that makes it effortless to organize and manage database tools. It provides built-in primitives—actual implementations of low-level functions like executing SQL—across different database systems (e.g., PostgreSQL, MySQL), ready to use without us having to code anything.
@@
In order to write any successful query over relational databases, the first step is always to understand the schema. In theory, relational databases should have table and column names that are unambiguous, column types should match their intended semantics (e.g., an age column is a number and not text), and keys and foreign keys should be explicitly declared in the schema. In practice, this is rarely the case. Corporate or institutional databases have dozens or even hundreds of tables, each with dozens of attributes, and the table or column names are frequently non-descriptive. For example even the relatively well organized Crime Data for Seattle (Seattle Police Department 2025a) has opaque column names like NIBRS Group AB or Beat.
@@
Once a reasonable understanding of the database’s schema has been established, the next step is to interact and play around with the data itself. For relational databases, this means writing SQL. The NL to SQL problem consists of automatically converting a natural language question over a relational database (with known schema) to an SQL query (Li and Jagadish 2014; Sen et al. 2020). Despite lots of progress in this space and the availability of popular benchmarks—like Spider (Yu et al. 2018), KaggleDBQA (Lee, Polozov, and Richardson 2021) and BIRD (Li et al. 2023)—NL2SQL remains a challenging problem for real-world scenarios, because of schema complexity, query ambiguity, and semantic mismatch (Floratou et al. 2024).
@@ 3 Verification Example
We set out to verify this claim using Thucy. We have access to a PostgreSQL database with the City of Seattle’s official crime data. We had downloaded this data from the Seattle Police Department (2025b) in the form of a CSV file, then uploaded it into PostgreSQL. It contains roughly 1.51.5M rows of all recorded crimes from 2008 to the present. To verify claims like that by the city’s attorney, we invoked Thucy with the verbatim claim above. After a few minutes, our multi-agent system produces the following report and verdict (excerpted verbatim):
@@
An important feature of Thucy is that it also returns the explanatory SQL query (or queries) that lead it to its verdict. These queries can be directly inspected, executed, and reproduced by expert users. In our example, Thucy generated the query shown in Figure 5. In essence, the query groups crimes by year and category, and then counts the number of incidents—exactly what we would expect for this verification. The output of the query is also shown in the figure. It was easy to run this query ourselves and confirm the correctness of Thucy’s verdict; we show the answers in the figure. We also checked these results on the interactive crime dashboard of the City of Seattle (Seattle Police Department 2025a), and got the same results. [^2] In the dashboard, make sure that “all” is selected in Precinct.
@@ 4 Experiments
We conduct experiments on TabFact (Chen et al. 2020), a widely used benchmark for fact verification over Wikipedia tables. The task is to determine whether a claim holds given the evidence in a relational table. The claim is labeled “false” if any part of it conflicts with the data from the table. Many cases involve subtle linguistic reasoning and common sense. Following all prior work (Nguyen et al. 2025; Zhu et al. 2024; Zhang et al. 2024), we evaluate on the small test split of TabFact, which contains roughly 2k fact-table pairs.
@@
We compare against recent fact-verification systems that all rely on LLMs, as these have achieved state-of-the-art performance (Zhu et al. 2024). We do not re-implement the baselines; instead, we report the results provided in their original papers for the same task and dataset.
We compare against BINDER (Cheng et al. 2023), DATER (Ye et al. 2023), CoTable (Wang et al. 2024), ReActTable (Zhang et al. 2024), AutoTQA (Zhu et al. 2024), and POS (Nguyen et al. 2025).
Among them, AutoTQA is particularly relevant, as it also builds a multi-agent system and is the only one in the literature to also support cross-table querying. Their agents follow a cyclic orchestration pattern—executing, critiquing, and refining plans in a loop. Our approach differs in two main ways: 1 Thucy is agnostic to the underlying data environment, and 2 it provides concrete traceable evidence alongside the answers. We also take a different stance on agent orchestration: instead of cyclic pattern, we employ decoupled, specialized expert-agents. This choice is validated by recent successful applications in industry (Anthropic 2025).
@@ Results
BINDER (Cheng et al. 2023) Codex 85.1%
BINDER (Nguyen et al. 2025) GPT-4o-mini 84.6%
DATER (Ye et al. 2023) Codex 85.6%
DATER (Nguyen et al. 2025) GPT-4o-mini 81.0%
CoTable (Wang et al. 2024) PaLM 2 86.6%
CoTable (Nguyen et al. 2025) GPT-4o-mini 84.2%
ReAcTable (Zhu et al. 2024) GPT-4 83.4%
ReAcTable (Zhu et al. 2024) GPT-4-turbo 85.0%
ReAcTable (Zhang et al. 2024) Codex 86.1%
AutoTQA (Zhu et al. 2024) GPT-4 87.4%
AutoTQA (Zhu et al. 2024) GPT-4-turbo 88.7%
POS (Nguyen et al. 2025) GPT-4o-mini 82.7%
@@ 5 A Journalistic Tale
A very recent article by MyNorthwest (2025) claimed that violent crime in downtown Seattle had “plummeted” during the summer months compared to the same period last year. The second sentence of the article reads:
@@
Within just a few hours, other outlets—including Kiro7 (2025) and Yahoo News (2025)—had picked up and republished the same story, all citing the original source.
Naturally, having built Thucy and with the crime dataset from the Seattle Police Department (2025b) already in hand, we were eager to see what it would say. We submitted the exact claim verbatim and waited a few minutes for the answer. Unlike the earlier statement by Seattle’s City Attorney (Sections 1 and 3), however, this one was trickier. The term, “downtown”, in particular, can mean different things depending on who you ask. After a few minutes, Thucy returned the following report (excerpted verbatim):
@@ Conclusion
After further investigation on the Web, we finally uncovered the source of the confusion. The original news source came from a different article, published by the Downtown Seattle Association (2025), which stated:
@@
With the extra M sector information at hand, Thucy was able to verify it. We noticed that this time Thucy took a different route: it filtered the data using attributes like sector. We also verified Thucy ’s findings by cross-checking the results on the interactive crime dashboard of the Seattle Police Department (2025a). The numbers match perfectly. [^4] When reproducing the results, after choosing the year and offense category, keep only beats M1–M3 selected; this is sector M.
Even though we are not journalists, this whole process convinced us even further of the urgent need for journalistic tools that actually produce the concrete SQL evidence of their verdict. Thucy is one of them. It doesn’t just give a verdict—the story doesn’t just end there. It can be transformed, magnified, and turned to something greater. This is what Cohen et al. (2011) envisioned long ago in their pioneering seminal work on computational journalism.
@@ 6 Limitations & Future Work
Coming back to the previous example, when we first gave Thucy the ambiguous claim about downtown Seattle, it made some assumptions about the neighborhood. It then filtered this attribute with the values it considered as downtown. So far so good—but what Thucy missed was that this column has many missing values. In fact, roughly 50% of them are missing. Since Thucy also returns the concrete SQL, we were able to spot this immediately.
While reviewing Thucy’s generated queries, we often ended up doing the opposite of what it was built for—trying to validate the original claims instead of just verifying them. More formally, we were searching for the missing predicates that would make a claim true. When Thucy flagged something as Inaccurate, we would dig into the data ourselves to see if it might have missed something. It could be the case that Thucy got tricked by missing values, misleading column names, or simply extra context buried somewhere in the database. Whatever the case, it makes us wonder: what if we built a multi-agent system that did exactly that? One that tries to support a claim and find evidence for it—a kind of “twin” system running in parallel to Thucy. Ideally, the two would agree on the final verdict. If not, that’s when things get interesting. What was missed? Are there consistent patterns? What can we do to avoid them?
Another direction we want to explore is controlling how much the agents rely on assumptions. Assumptions can be useful, but they can also introduce subtle errors (for example, using the wrong current date). We want to experiment with ways to make these assumptions explicit or better grounded. One idea is to create another specialized expert-agent that searches the web (e.g., Google).
Lastly, multi-agent systems like Thucy burn through tokens fast (Anthropic 2025). In our case, fact-checking all 4K examples in our experiments cost about \$183.9 in total, mainly because Thucy must explore the underlying data environment from scratch every single time. That comes out to roughly 5¢ per example. In messy real-world fact-checking scenarios like the one discussed earlier (Section 5), the cost rises to 20¢ per verification.
Still, we believe that our journalistic use case is high-stakes enough that this trade-off is worthwhile. After all, we can easily imagine journalists at the New York Times being more than happy to spend a few dollars to have their articles stamped by Thucy as Verified and fault-proof.
@@ 7 Conclusion
Building Thucy has been a continuous learning experience. The field of multi-agent systems is evolving at an astronomical pace, with new frameworks, research breakthroughs, and reports emerging almost weekly. We hope that the insights and design choices shared in this paper will help others accelerate their work and keep the momentum going.
@@ 8 Acknowledgments
This work was supported by NSF III 2507117, NSF SHF 2312195, and NSF IIS 2314527.

### v2

@@ 1 Introduction
In the Annual Report released last year by the SeattleCityAnnualReport_2024, we read the following:
@@
It turns out, the City of Seattle publicly provides an official crime dataset (SeattleCrimeDataset)—with all crimes from 2008 until now—that is structured, detailed, and updated almost daily. In principle, that is all you would need to verify such a claim. In practice, though, very few people ever try. Most will simply take the statement at face value and move on, keeping the comforting thought that “Seattle is safer now” somewhere in the back of their mind.
@@
We draw inspiration from the work of Thucy(dides), the Athenian historian (460–400 BC) who wrote the History of the Peloponnesian War between Sparta and Athens. “Thucydides has been dubbed the father of scientific history by those who accept his claims to have applied strict standards of impartiality and evidence-gathering” (wiki:Thycidides).
@@ 2 Architecture
BINDER (DBLP:conf/iclr/ChengX0LNHXROZS23)
DATER (DBLP:conf/sigir/YeHYLHL23)
CoTable (DBLP:conf/iclr/0002ZLEP0MFSLP24)
ReAcTable (DBLP:journals/pvldb/ZhangHFCDP24)
AutoTQA (DBLP:journals/pvldb/ZhuCXLSZSTL24)
POS (DBLP:journals/tmlr/NguyenBSKNL25)
@@
To enable this flexibility, we must address a fundamental challenge: LLMs, no matter how capable, are inherently disconnected from external data sources—they can only operate in isolation, with no way to interact with databases. Agents bridge this gap by using tools. A tool acts as an interface to external capabilities, allowing agents to interact with, perceive, and affect their environment. In general, tools can include capabilities that perform mathematical calculations, or read files from disk, or query a database. Each agent has a fixed collection of such tools. At runtime, the agent [^1] More precisely, it is the LLM that makes this decision, though we often use “agent” and “LLM” interchangeably in such contexts autonomously decides which tool to invoke, how to call it, and when to use it; the tool’s output is then fed back into the reasoning loop (DBLP:conf/iclr/YaoZYDSN023). This interactive feedback cycle between reasoning, action, and observations forms the backbone of modern agentic AI.
@@
The Model Context Protocol (MCP), introduced by AnthropicMCP, standardizes how AI applications connect to different data sources, effectively eliminating the need for custom connections for each new AI model and external system allowing us to direct our energy elsewhere—away from repetitive boilerplate code. MCP simplifies and streamlines the tool building process, and has already been adopted by industry (PaypalMCProllout; DatabricksMCProllout; AzureMCProllout; SnowflakeMCProllout).
@@
We use Google’s MCP Toolbox for Databases (BlogPostGoogleMCP), a framework that makes it effortless to organize and manage database tools. It provides built-in primitives—actual implementations of low-level functions like executing SQL—across different database systems (e.g., PostgreSQL, MySQL), ready to use without us having to code anything.
@@
In order to write any successful query over relational databases, the first step is always to understand the schema. In theory, relational databases should have table and column names that are unambiguous, column types should match their intended semantics (e.g., an age column is a number and not text), and keys and foreign keys should be explicitly declared in the schema. In practice, this is rarely the case. Corporate or institutional databases have dozens or even hundreds of tables, each with dozens of attributes, and the table or column names are frequently non-descriptive. For example even the relatively well organized Crime Data for Seattle (SeattleCrimeDashboard2025) has opaque column names like NIBRS Group AB or Beat.
@@
Once a reasonable understanding of the database’s schema has been established, the next step is to interact and play around with the data itself. For relational databases, this means writing SQL. The NL to SQL problem consists of automatically converting a natural language question over a relational database (with known schema) to an SQL query (DBLP:conf/sigmod/LiJ14; DBLP:journals/pvldb/SenLQOEDSMSS20). Despite lots of progress in this space and the availability of popular benchmarks—like Spider (DBLP:conf/emnlp/YuZYYWLMLYRZR18), KaggleDBQA (DBLP:conf/acl/LeePR20) and BIRD (DBLP:conf/nips/LiHQYLLWQGHZ0LC23)—NL2SQL remains a challenging problem for real-world scenarios, because of schema complexity, query ambiguity, and semantic mismatch (DBLP:conf/cidr/FloratouPZDHTCA24).
@@ 3 Verification Example
We set out to verify this claim using Thucy. We have access to a PostgreSQL database with the City of Seattle’s official crime data. We had downloaded this data from the SeattleCrimeDataset in the form of a CSV file, then uploaded it into PostgreSQL. It contains roughly 1.51.5M rows of all recorded crimes from 2008 to the present. To verify claims like that by the city’s attorney, we invoked Thucy with the verbatim claim above. After a few minutes, our multi-agent system produces the following report and verdict (excerpted verbatim):
@@
An important feature of Thucy is that it also returns the explanatory SQL query (or queries) that lead it to its verdict. These queries can be directly inspected, executed, and reproduced by expert users. In our example, Thucy generated the query shown in Figure 5. In essence, the query groups crimes by year and category, and then counts the number of incidents—exactly what we would expect for this verification. The output of the query is also shown in the figure. It was easy to run this query ourselves and confirm the correctness of Thucy’s verdict; we show the answers in the figure. We also checked these results on the interactive crime dashboard of the City of Seattle (SeattleCrimeDashboard2025), and got the same results. [^2] In the dashboard, make sure that “all” is selected in Precinct.
@@ 4 Experiments
We conduct experiments on TabFact (DBLP:conf/iclr/ChenWCZWLZW20), a widely used benchmark for fact verification over Wikipedia tables. The task is to determine whether a claim holds given the evidence in a relational table. The claim is labeled “false” if any part of it conflicts with the data from the table. Many cases involve subtle linguistic reasoning and common sense. Following all prior work (DBLP:journals/tmlr/NguyenBSKNL25; DBLP:journals/pvldb/ZhuCXLSZSTL24; DBLP:journals/pvldb/ZhangHFCDP24), we evaluate on the small test split of TabFact, which contains roughly 2k fact-table pairs.
@@
We compare against recent fact-verification systems that all rely on LLMs, as these have achieved state-of-the-art performance (DBLP:journals/pvldb/ZhuCXLSZSTL24). We do not re-implement the baselines; instead, we report the results provided in their original papers for the same task and dataset.
We compare against BINDER (DBLP:conf/iclr/ChengX0LNHXROZS23), DATER (DBLP:conf/sigir/YeHYLHL23), CoTable (DBLP:conf/iclr/0002ZLEP0MFSLP24), ReActTable (DBLP:journals/pvldb/ZhangHFCDP24), AutoTQA (DBLP:journals/pvldb/ZhuCXLSZSTL24), and POS (DBLP:journals/tmlr/NguyenBSKNL25).
Among them, AutoTQA is particularly relevant, as it also builds a multi-agent system and is the only one in the literature to also support cross-table querying. Their agents follow a cyclic orchestration pattern—executing, critiquing, and refining plans in a loop. Our approach differs in two main ways: 1 Thucy is agnostic to the underlying data environment, and 2 it provides concrete traceable evidence alongside the answers. We also take a different stance on agent orchestration: instead of cyclic pattern, we employ decoupled, specialized expert-agents. This choice is validated by recent successful applications in industry (AnthropicMultiAgentSystem).
@@ Results
BINDER (DBLP:conf/iclr/ChengX0LNHXROZS23) Codex 85.1%
BINDER (DBLP:journals/tmlr/NguyenBSKNL25) GPT-4o-mini 84.6%
DATER (DBLP:conf/sigir/YeHYLHL23) Codex 85.6%
DATER (DBLP:journals/tmlr/NguyenBSKNL25) GPT-4o-mini 81.0%
CoTable (DBLP:conf/iclr/0002ZLEP0MFSLP24) PaLM 2 86.6%
CoTable (DBLP:journals/tmlr/NguyenBSKNL25) GPT-4o-mini 84.2%
ReAcTable (DBLP:journals/pvldb/ZhuCXLSZSTL24) GPT-4 83.4%
ReAcTable (DBLP:journals/pvldb/ZhuCXLSZSTL24) GPT-4-turbo 85.0%
ReAcTable (DBLP:journals/pvldb/ZhangHFCDP24) Codex 86.1%
AutoTQA (DBLP:journals/pvldb/ZhuCXLSZSTL24) GPT-4 87.4%
AutoTQA (DBLP:journals/pvldb/ZhuCXLSZSTL24) GPT-4-turbo 88.7%
POS (DBLP:journals/tmlr/NguyenBSKNL25) GPT-4o-mini 82.7%
@@ 5 A Journalistic Tale
A very recent article by MyNorthwestClaimOct2025 claimed that violent crime in downtown Seattle had “plummeted” during the summer months compared to the same period last year. The second sentence of the article reads:
@@
Within just a few hours, other outlets—including Kiro7ClaimOct2025 and YahooNewsClaimOct2025—had picked up and republished the same story, all citing the original source.
Naturally, having built Thucy and with the crime dataset from the SeattleCrimeDataset already in hand, we were eager to see what it would say. We submitted the exact claim verbatim and waited a few minutes for the answer. Unlike the earlier statement by Seattle’s City Attorney (Sections 1 and 3), however, this one was trickier. The term, “downtown”, in particular, can mean different things depending on who you ask. After a few minutes, Thucy returned the following report (excerpted verbatim):
@@ Conclusion
After further investigation on the Web, we finally uncovered the source of the confusion. The original news source came from a different article, published by the OriginalClaimOct2025, which stated:
@@
With the extra M sector information at hand, Thucy was able to verify it. We noticed that this time Thucy took a different route: it filtered the data using attributes like sector. We also verified Thucy ’s findings by cross-checking the results on the interactive crime dashboard of the SeattleCrimeDashboard2025. The numbers match perfectly. [^4] When reproducing the results, after choosing the year and offense category, keep only beats M1–M3 selected; this is sector M.
Even though we are not journalists, this whole process convinced us even further of the urgent need for journalistic tools that actually produce the concrete SQL evidence of their verdict. Thucy is one of them. It doesn’t just give a verdict—the story doesn’t just end there. It can be transformed, magnified, and turned to something greater. This is what DBLP:conf/cidr/CohenLYY11 envisioned long ago in their pioneering seminal work on computational journalism.
@@ 6 Limitations & Future Work
Dirty Data. Coming back to the previous example, when we first gave Thucy the ambiguous claim about downtown Seattle, it made some assumptions about the neighborhood. It then filtered this attribute with the values it considered as downtown. So far so good—but what Thucy missed was that this column has many missing values. In fact, roughly 50% of them are missing. Since Thucy also returns the concrete SQL, we were able to spot this immediately.
Assumptions & Ambiguity. Another direction we want to explore is controlling how much the agents rely on assumptions. Assumptions are useful as this is the only way to combat ambiguity in both the data and user questions. However, they can also introduce subtle errors (for example, using the wrong current date). We want to experiment with ways to make these assumptions better grounded. One idea is to create another specialized expert-agent that searches the web.
Quantitative Evaluation. In the evaluation of Thucy, we used TabFact (DBLP:conf/iclr/ChenWCZWLZW20). There is, however, a mismatch: we propose a system that can navigate data environments with many databases and tables, while our evaluation is conducted on a single-table benchmark. This is indeed the case, but to the best of our knowledge there is no fact-verification benchmark in the literature that focuses on large-scale cross-table or cross-database data. We believe this is an important next step for fact-verification, and we are actively working on creating one.
Ablation Studies. As we were building Thucy, we manually tested and refined each agent, observing both their individual behavior and their interactions within the full system. However, in this work we do not present systematic ablation studies. A careful evaluation and a systematic study of the contribution of each component—both in isolation and by removing individual agents from the system—is warranted.
Stateless Expert-Agents. Each expert agent does not preserve memory across tasks. This is a deliberate design choice, as it allows Thucy to operate in dynamic data environments. However, this increases cost, since agents must re-discover information across tasks.
Expensive. Lastly, multi-agent systems like Thucy burn through tokens fast (AnthropicMultiAgentSystem). In our case, fact-checking all 4K examples in our experiments cost about \$183.9 in total. That comes out to roughly 5¢ per example. In messy real-world fact-checking scenarios like the one discussed earlier (Section 5), the cost rises to 20¢ per verification. Still, we believe that our journalistic use case is high-stakes enough that this trade-off is worthwhile. After all, we can easily imagine journalists at the New York Times being more than happy to spend a few dollars to have their articles stamped by Thucy as verified and fault-proof.
@@ 8 Acknowledgments
We thank the reviewers. This work was supported by NSF III 2507117, NSF SHF 2312195, and NSF IIS 2314527.